a.   If we immerse a circular coil of 190 mm diameter into a material of relative permeability of 3 and the coil conducts a current of 8A, find the magnetic induction at the center of the coil.

b.   A solenoid with length 65 cm and radius 2 cm has 300 turns of copper wire around a non-magnetic material. If the current flowing through the wire is 3.2A, find the magnetic induction along the center of the solenoid.



1
Expert's answer
2021-11-21T17:27:19-0500

a)

"B=\frac{\mu\mu_0 I}{D}=1.6\cdot 10^{-4}~T,"

b)

"B=\mu_0I\frac N{2\pi r}=9.6\cdot 10^{-4}~T."